OpenStudy (anonymous):

Need Help to solve out this question. Find dy/dx: x= radical of tan (y^2).

zepdrix (zepdrix):

Was it too hard to understand when we did it earlier? :c

OpenStudy (anonymous):

that was wrong actually

zepdrix (zepdrix):

It wasn't wrong. I even checked it on wolfram. If you're not suppose to solve it implicitly, then you'll have an answer involving arctangents and such. It will look much different than what we came up with, but still correct.

OpenStudy (anonymous):

hmm but the answer I have from professor is \[(\sqrt{\tan(y)}/(y \sec ^{2}y)\]

OpenStudy (anonymous):

so i am not really sure for how to get to this answer, I am really stuck my brain is not working these problems are very hard for me I have 4 more which I really need help with.

zepdrix (zepdrix):

That's the exact answer I put in the post we were working on :( Did you just skip to the bottom i guess...?

OpenStudy (anonymous):

I am sorry Zepdrix it was my fault, I am really confuse

OpenStudy (anonymous):

This is the first time I am doing homework for Calculus and I am having tough time.

zepdrix (zepdrix):

Oh your y's look a little different :D hmm. Yah it's a really tough problem :C teacher is being mean lol

OpenStudy (anonymous):

lolz

OpenStudy (anonymous):

do you mind if you can teach me that problem again ??

zepdrix (zepdrix):

So which part are you getting stuck on? The problem gets really really messy later on. But do you understand how to start the problem? Taking the derivative of the `outermost function`.

OpenStudy (anonymous):

I am really sorry but I need to admit this that I really dont understand frm beginning

zepdrix (zepdrix):

lol XD

OpenStudy (anonymous):

:)

zepdrix (zepdrix):

We're differentiating this `implicitly`, meaning that we're NOT solving the function explicitly for \(y\) before taking a derivative. See the big pile of junk attached to the \(y\)? It makes it a lot different than problems that you've maybe dealt with in the past. Let's try an easier problem really quick.

OpenStudy (anonymous):

ok :)

zepdrix (zepdrix):

\[\large x=\left(y^3+1\right)^2\]Taking the derivative with respect to `x` gives us,\[\large 1=2(y^3+1)^1 \quad \color{royalblue}{(y^3+1)'}\] Do you understand why we get a 1 on the left? :O

OpenStudy (anonymous):

because if we take derivative of x it will give us 1

zepdrix (zepdrix):

mhm c: true story.

OpenStudy (anonymous):

yesss finally I got something right :D

zepdrix (zepdrix):

On the right side, we start by taking the derivative of the `outermost function`, which in this case happens to be the square on the outside. So we apply the power rule to that part, that's how the 2 comes down.

OpenStudy (anonymous):

ahan yup

zepdrix (zepdrix):

Then the chain rule tells us to do this.